The weights of students in a certain school are normally distributed with a mean weight of 66 kg. 10% have a weight greater than 70kg. What percentage of students weighs between 62kg and 66kg?


1
Expert's answer
2022-04-19T01:07:00-0400

The normal distribution is symmetric:

"P(W\le66)=P(W\ge66)=50\%\\
P(66-4=62\le W\le66)=\\
=P(66\le W\le70=66+4)=\\
=50\%-P(70\le W)=40\%"